# OpenGVL - Benchmarking Visual Temporal Progress for Data Curation

### v1

@@ Abstract
Data scarcity remains one of the most limiting factors in driving progress in robotics. However, the amount of available robotics data in the wild is growing exponentially, creating new opportunities for large-scale data utilization. Reliable temporal task completion prediction could help automatically annotate and curate this data at scale. The Generative Value Learning (GVL) approach was recently proposed, leveraging the knowledge embedded in vision-language models (VLMs) to predict task progress from visual observations. Building upon GVL, we propose OpenGVL, a comprehensive benchmark for estimating task progress across diverse challenging manipulation tasks involving both robotic and human embodiments. We evaluate the capabilities of publicly available open-source foundation models, showing that open-source model families significantly underperform closed-source counterparts, achieving only approximately of their performance on temporal progress prediction tasks. Furthermore, we demonstrate how OpenGVL can serve as a practical tool for automated data curation and filtering, enabling efficient quality assessment of large-scale robotics datasets. We release the benchmark along with the complete codebase at OpenGVL.
@@ 4 Data curation in the wild
To demonstrate the potential of OpenGVL for data curation, we analyzed various datasets recently shared on the Hugging Face LeRobot datasets hub. With over 13,000 datasets already published, automatic data curation and filtering have become essential for leveraging these datasets during the pre-training phase. We show how OpenGVL can easily identify different dataset issues ranging from unclear task definitions and instructions to occluded sensors and failed/out-of-distribution examples that can disrupt training, as observed previously [20].

### v2

@@ Abstract
Data scarcity remains one of the most limiting factors in driving progress in robotics. However, the amount of available robotics data in the wild is growing exponentially, creating new opportunities for large-scale data utilization. Reliable temporal task completion prediction could help automatically annotate and curate this data at scale. The Generative Value Learning (GVL) approach was recently proposed, leveraging the knowledge embedded in vision-language models (VLMs) to predict task progress from visual observations. Building upon GVL, we propose OpenGVL, a comprehensive benchmark for estimating task progress across diverse challenging manipulation tasks involving both robotic and human embodiments. We evaluate the capabilities of publicly available open-source foundation models, showing that open-source model families significantly underperform closed-source counterparts, achieving only approximately of their performance on temporal progress prediction tasks. Furthermore, we demonstrate how OpenGVL can serve as a practical tool for automated data curation and filtering, enabling efficient quality assessment of large-scale robotics datasets. We release the benchmark along with the complete codebase.
@@ 4 Data curation in the wild
To demonstrate the potential of OpenGVL for data curation, we analyzed various datasets recently shared on the Hugging Face LeRobot datasets hub. With over 13,000 datasets already published, automatic data curation and filtering have become essential for leveraging these datasets during the pre-training phase.
We show how OpenGVL can easily identify different dataset issues ranging from unclear task definitions and instructions to occluded sensors and failed/out-of-distribution examples that can disrupt training, as observed previously [20].
@@ Appendix A Prompt
Appendix B Detailed evaluation results
| Model | Dataset | Ctx | VOC Mean | VOC Std | Std Err | Mism. | Empty |
| gemini-2.5-flash-lite | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| gemini-2.5-pro | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| google/gemma-3-12b-it | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| Model | Dataset | Ctx | VOC Mean | VOC Std | Std Err | Mism. | Empty |
| google/gemma-3-27b-it | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| google/gemma-3-4b-it | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| moonshotai/Kimi-VL-A3B-Thinking-2506 | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| Model | Dataset | Ctx | VOC Mean | VOC Std | Std Err | Mism. | Empty |
| nvidia/Cosmos-Reason1-7B | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| Qwen/Qwen2.5-VL-32B-Instruct | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| Qwen/Qwen2.5-VL-3B-Instruct | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| Model | Dataset | Ctx | VOC Mean | VOC Std | Std Err | Mism. | Empty |
| Qwen/Qwen2.5-VL-7B-Instruct | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| XiaomiMiMo/MiMo-VL-7B-RL-2508 | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||
| zai-org/GLM-4.1V-9B-Thinking | lerobot/berkeley_mvp | ||||||
| lerobot/cmu_stretch | |||||||
| lerobot/nyu_door_opening_surprising_effectiveness | |||||||
| lerobot/nyu_franka_play_dataset | |||||||
| Hidden Task 1 | |||||||
| Hidden Task 2 | |||||||